### v2

# Duality for Fitting’s Multi-valued Modal logic via bitopology and bi-Vietoris coalgebra

Abstract
Fitting’s Heyting-valued logic and Heyting-valued modal logic have already been studied from an algebraic viewpoint. In addition to algebraic axiomatizations with the completeness of Fitting’s Heyting-valued logic and Heyting-valued modal logic, both topological and coalgebraic dualities have also been developed for algebras of Fitting’s Heyting-valued modal logic. Bitopological methods have recently been employed to investigate duality for Fitting’s Heyting-valued logic. However, the concepts of bitopology and biVietoris coalgebras are conspicuously absent from the development of dualities for Fitting’s many-valued modal logic. With this study, we try to bridge that gap. We develop a bitopological duality for algebras of Fitting’s Heyting-valued modal logic. We construct a bi-Vietoris functor on the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued ($\mathcal{L}$ is a Heyting algebra) pairwise Boolean spaces. Finally, we obtain a dual equivalence between categories of biVietoris coalgebras and algebras of Fitting’s Heyting-valued modal logic. As a result, we conclude that Fitting’s many-valued modal logic is sound and complete with respect to the coalgebras of a biVietoris functor. We discuss the application of this coalgebraic approach to bitopological duality.
keywords:
Bitopology, Fitting’s modal logic, Duality, Coalgebra
1 Introduction
Algebraic axiomatization of a modified version of Fitting’s Heyting-valued modal logic has already been addressed in [11]. Maruyama [12] proposed Jónsson-Tarski topological duality (see [17, 18, 19]) for $\mathcal{L}$$-\mathcal{ML}$-algebras (algebras of Fitting’s Heyting-valued modal logic). Jónsson-Tarski duality for $\mathcal{L}$$-\mathcal{ML}$-algebras is essentially a $\mathcal{L}$-valued version of Jónsson-Tarski duality for modal algebras.
We aim to construct a bitopological duality for algebras of Fitting’s Heyting- valued modal logic by setting up a notion of $PRBS_{\mathcal{L}}$ as a category of $\mathcal{L}$-valued pairwise Boolean spaces with a binary relation. As a result, natural duality theory for modal algebras is extended in the context of bitopological languages. The main results are bitopological and coalgebraic dualities for $\mathcal{L}$$-\mathcal{ML}$-algebras, where $\mathcal{L}$ is a semi-primal algebra having a bounded lattice reduct. Our general theory extends both the Jónsson-Tarski duality and the Abramsky-Kupke-Kurz-Venema coalgebraic duality [1, 20] in the setting of bitopological language. Furthermore, it introduces a novel coalgebraic duality for $\mathcal{L}$$-\mathcal{ML}$-algebras.
An exemplary story in coalgebraic logic can be found in [21]. The Stone duality [22] between Boolean algebras and sets represents the syntax and semantics of a propositional logic. The algebras and coalgebras of the endofunctors define the syntax and semantics of the modal propositional logic. As an illustration, the modal logic $K$ and Kripke semantics derive from the Stone duality by taking an endofunctor on sets. So, in acceptable circumstances, we can achieve duality between the relevant algebras and coalgebras. In addition to demonstrating the fact that the widely recognised Stone duality could be articulated in coalgebraic terms, Abramsky [1] discovered a coalgebraic approach to the Jónsson-Tarski duality for modal algebras (for more information, see also [20]). It is specifically noted that the category of Boolean spaces is isomorphic to the category of descriptive general Kripke frames. Esakia [23] also noticed this connection. Coalgebras for the Vietoris functor on the category of Boolean spaces , also known as Stone spaces, can thus be used to represent semantics for modal logic. In [24], the author showed that coalgebras of a Vietoris functor on the category of Priestley spaces, i.e., compact, totally ordered disconnected spaces, provide sound and complete semantics for positive modal logic. The author in [7] subsequently developed a Vietoris functor on the category of pairwise Stone spaces, whereby coalgebras for that defined functor yield sound and complete semantics for the positive modal logic.
In this study, we seek to demonstrate how coalgebras for the biVietoris functor on the category $PBS_{\mathcal{L}}$ of $\mathcal{L}$-valued version of pairwise Boolean spaces and pairwise continuous maps can be used to understand the semantics of Fitting’s many-valued modal logic.
Our study develops both bitopological and coalgebraic dualities for algebras of Fitting’s many-valued modal logic, building upon previous research [14, 13, 7, 8]. Maruyama et al. [13] developed both topological and coalgebraic dualities for general algebraic structures encompassing algebras of Fitting’s many-valued modal logic. Our research directly contributes to the results of the publication [13]; specifically, we present bitopological versions of those dualities for Fitting’s many-valued modal logic.
The structure of the paper is as follows:
In Section 2, we review the fundamental notions of bitopological spaces and algebras of Fitting’s Heyting-valued logic, denoted as $\mathcal{L}$-$\mathcal{VL}$-algebras, and we also discuss the bitopological duality theorem for $\mathcal{L}$-$\mathcal{VL}$-algebras. Section 3 contains our first result, i.e., bitopological duality for Fitting’s Heyting-valued modal logic. In Section 4, we introduce the concept of pairwise Vietoris spaces and construct an endofunctor $V_{\mathcal{L}}^{bi}$ on the category $PBS_{\mathcal{L}}$. Finally, we show the coalgebraic duality for Fitting’s Heyting-valued modal logic. In Section 5, we conclude the paper by outlining several potential future research directions.
2 Preliminaries
The readers are believed to be conversant with the fundamental ideas of category theory and topology. For details on universal algebra and lattice theory, we refer the reader to [2, 3]. To get more information about category theory, see [4].
2.1 Bitopological spaces
A triple $(X,\tau_{1},\tau_{2})$, where $(X,\tau_{1})$ and $(X,\tau_{2})$ are topological spaces, is called a bitopological space. Consider $\delta_{1}$ and $\delta_{2}$ represent, respectively, the collections of $\tau_{1}$-closed sets and $\tau_{2}$-closed sets. We set $\beta_{1}=\tau_{1}\cap\delta_{2}$ and $\beta_{2}=\tau_{2}\cap\delta_{1}$.
Definition 1 ([6]).
  1. (i)
    A bitopoological space $(X,\tau_{1},\tau_{2})$ is said to be pairwise Hausdorff space if for every pair $(x,y)$ of distinct points $x,y\in X$ there exists disjoint open sets $U_{x}\in\tau_{1}$ and $U_{y}\in\tau_{2}$ containing $x$ and $y$, respectively.
  2. (ii)
    A bitopoological space $(X,\tau_{1},\tau_{2})$ is said to be pairwise zero-dimensional if $\beta_{1}$ is a basis for $\tau_{1}$ and $\beta_{2}$ is a basis for $\tau_{2}$.
  3. (iii)
    A bitopoological space $(X,\tau_{1},\tau_{2})$ is said to be pairwise compact if the topological space $(X,\tau)$, where $\tau=\tau_{1}\vee\tau_{2}$, is compact.
According to Alexander’s Lemma( a classical result in general topology), the idea of pairwise compactness described in Definition 1 is equivalent to the condition that every cover $\{U:U\in\tau_{1}\cup\tau_{2}\}$ of X has a finite subcover. A pairwise Boolean space is a bitopological space that is pairwise Hausdorff, pairwise zero-dimensional, and pairwise compact. A map $f:(P,\tau_{1},\tau_{2})\to(P_{1},\tau_{1}^{1},\tau_{2}^{1})$ is said to be pairwise continuous if the map $f:(P,\tau_{i})\to(P_{1},\tau_{i}^{1})$ is continuous for $i\in\{1,2\}$. Pairwise Boolean spaces and pairwise continuous maps form a category, denoted by $PBS$.
Proposition 1 ([7]).
If $T_{1}$ and $T_{2}$ are subbasis for the topologies $\tau_{1}$ and $\tau_{2}$, respectively, then $T_{1}\cup T_{2}$ is a subbasis for the topology $\tau_{1}\vee\tau_{2}$.
Proposition 2 ([7]).
Let $(X,\tau_{1},\tau_{2})$ be a pairwise compact bitopological space. Consider a finite collection $\{C_{i}:C_{i}\in\delta_{1}\cup\delta_{2},i=1,2,\cdots,n\}$ of subsets of $X$. Then $\bigcap_{i=1}^{n}C_{i}$ is pairwise compact.
It is clear from the above proposition that any $\tau_{1}$-closed or $\tau_{2}$-closed subset of a pairwise compact space $X$ is pairwise compact.
2.2 Fitting’s Heyting-valued logics
Fitting [10] proposed $\mathcal{L}$-valued logics and $\mathcal{L}$-valued modal logics for a finite distributive lattice $\mathcal{L}$(i.e., $\mathcal{L}$ is a Heyting algebra) in 1991. Maruyama [11] introduced algebraic axiomatization of Fitting’s logics. In [11] the author studied Fitting’s Heyting-valued logic and Heyting-valued modal logic without regard for fuzzy truth constants other than 00 and $1$, and added a new operation $T_{\ell}(-)$, $\ell\in\mathcal{L}$. From a logical perspective, $T_{\ell}(p)$ infers that the truth value of a proposition $p$ is $\ell$. The operations of $\mathcal{L}$-valued logic, denoted by $\mathcal{L}$-$\mathcal{VL}$, are $\vee,\wedge,\rightarrow,0,1$ and $T_{\ell}(-)$, $\ell\in\mathcal{L}$, where ,$\vee$,$\wedge$,$\rightarrow$ , ∧ , → are binary operations, 00 and $1$ are nullary operations and $T_{\ell}$ is a unary operation. For $\ell_{1},\ell_{2}\in\mathcal{L}$, $\ell_{1}\rightarrow\ell_{2}$ means the pseudo-complement of $\ell_{1}$ relative to $\ell_{2}$.
The following lemmas describe some term functions.
Lemma 1.
Define a function $f:\mathcal{L}^{4}\to\mathcal{L}$ by
$f(\ell_{1},\ell_{2},\ell_{3},\ell_{4})=\left\{\begin{array}[]{ll}\ell_{3}&(\ell_{1}=\ell_{2})\\ \ell_{4}&(\ell_{1}\neq\ell_{2})\\ \end{array}\right.$
Then, $f$ is a term function of $\mathcal{L}$.
Lemma 2.
For every $\ell\in\mathcal{L}$, define $T_{\ell}:\mathcal{L}\to\mathcal{L}$ by
$T_{\ell}(\ell^{\prime})=\left\{\begin{array}[]{ll}1&(\ell^{\prime}=\ell)\\ 0&(\ell^{\prime}\neq\ell)\\ \end{array}\right.$
Then, $T_{\ell}$ is a term function of $\mathcal{L}$.
Lemma 3.
Let $\ell\in\mathcal{L}$. Then the function $U_{\ell}:\mathcal{L}\to\mathcal{L}$ defined by
$U_{\ell}(\ell^{\prime})=\left\{\begin{array}[]{ll}1&(\ell^{\prime}\geq\ell)\\ 0&(\ell^{\prime}\ngeq\ell)\\ \end{array}\right.$
, is a term function of $\mathcal{L}$.
We recall the idea of $\mathcal{L}$-$\mathcal{VL}$-algebras, which provides sound and complete semantics of $\mathcal{L}$-valued logic $\mathcal{L}$-$\mathcal{VL}$.
Definition 2 ([11]).
An algebraic structure $(\mathcal{A},\wedge,\vee,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$ is said to be a $\mathcal{L}$-$\mathcal{VL}$-algebra iff the following conditions hold for any $\ell_{1},\ell_{2}\in\mathcal{L}$, and $a,b\in\mathcal{A}$:
  1. (i)
    $(\mathcal{A},\wedge,\vee,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$ is a Heyting algebra;
  2. (ii)
    $T_{\ell_{1}}(a)\wedge T_{\ell_{2}}(b)\leq T_{\ell_{1}\rightarrow\ell_{2}}(a\rightarrow b)\wedge T_{\ell_{1}\wedge\ell_{2}}(a\wedge b)\wedge T_{\ell_{1}\vee\ell_{2}}(a\vee b)$;
    $T_{\ell_{2}}(a)\leq T_{T_{\ell_{1}}(\ell_{2})}(T_{\ell_{1}}(a))$;
  3. (iii)
    $T_{0}(0)=1$; $T_{\ell}(0)=0$ ($\ell\neq 0$); $T_{1}(1)=1$; $T_{\ell}(1)=0$, if $\ell\neq 1$;
  4. (iv)
    $\bigvee\{T_{\ell}(a):\ell\in\mathcal{L}\}=1$; $T_{\ell_{1}}(a)\vee(T_{\ell_{2}}(a)\rightarrow 0)=1$;
    $T_{\ell_{1}}(a)\wedge T_{\ell_{2}}(a)=0$, $(\ell_{1}\neq\ell_{2})$;
  5. (v)
    $T_{1}(T_{\ell}(a))=T_{\ell}(a)$, $T_{0}(T_{\ell}(a))=T_{\ell}(a)\rightarrow 0$, $T_{\ell_{2}}(T_{\ell_{1}}(a))=0$, $(\ell_{2}\neq 0,1)$;
  6. (vi)
    $T_{1}(a)\leq a$, $T_{1}(a\wedge b)=T_{1}(a)\wedge T_{1}(b)$;
  7. (vii)
    $\bigwedge_{\ell\in\mathcal{L}}(T_{\ell}(a)\leftrightarrow T_{\ell}(b))\leq(a\leftrightarrow b)$.
Note 1.
The class of all $\mathcal{L}$-$\mathcal{VL}$-algebras forms a variety (in the sense of universal algebra). If =$\mathcal{L}=\{0,1\}$, then $\mathcal{L}$-$\mathcal{VL}$-algebras become Boolean algebras.
Definition 3.
A function between $\mathcal{L}$-$\mathcal{VL}$-algebras is said to be homomorphism if it preserves the operations $\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1$.
Let $\mathcal{VA}_{\mathcal{L}}$ denote the category of $\mathcal{L}$-$\mathcal{VL}$-algebras.
$\mathcal{L}$-valued modal logic denoted by $\mathcal{L}$$-\mathcal{ML}$, is defined by $\mathcal{L}$-valued Kripke semantics. The idea of $\mathcal{L}$-valued Kripke semantics can be found in [12]. The operations of $\mathcal{L}$-valued modal logic $\mathcal{L}$$-\mathcal{ML}$ are the operations of $\mathcal{L}$-$\mathcal{VL}$ and a unary operation $\Box$, called modal operation. We now recall the concept of $\mathcal{L}$$-\mathcal{ML}$-algebras, which define a sound and complete semantics for $\mathcal{L}$$-\mathcal{ML}$.
Definition 4 ([11]).
An algebraic structure $(\mathcal{A},\wedge,\vee,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),\Box,0,1)$ is said to be a $\mathcal{L}$$-\mathcal{ML}$-algebra iff it satisfies the following conditions:
  1. (i)
    $(\mathcal{A},\wedge,\vee,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$ is a $\mathcal{L}$-$\mathcal{VL}$-algebra;
  2. (ii)
    $\Box(a\wedge b)=\Box a\wedge\Box b$;
  3. (iii)
    $\Box U_{\ell}(a)=U_{\ell}(\Box a)$, $\forall\ell\in\mathcal{L}$, where the unary operation $U_{\ell}(\ell\in\mathcal{L})$ is defined by $U_{\ell}(a)=\bigvee\{T_{\ell^{\prime}}(a):\ell\leq\ell^{\prime},\ell^{\prime}\in\mathcal{L}\}$, $a\in\mathcal{A}$. Logically, it means that the truth value of $a$ is greater than or equal to $\ell$.
A homomorphism of $\mathcal{L}$$-\mathcal{ML}$-algebras is a function that preserves all the operations of $\mathcal{L}$-$\mathcal{VL}$-algebras and the modal operation $\Box$. Let $\mathcal{MA}_{\mathcal{L}}$ denote the category of $\mathcal{L}$$-\mathcal{ML}$-algebras and homomorphisms of $\mathcal{L}$$-\mathcal{ML}$-algebras.
For a Kripke frame $(P,\mathcal{R})$, $\mathcal{R}[x]=\{y\in P:x\mathcal{R}y\}$, where $x\in P$, and $\mathcal{R}^{-1}[P^{\prime}]=\{y\in P:\exists x\in P^{\prime},y\mathcal{R}x\}$, where $P^{\prime}\subseteq P$. We recall a modal operation $\Box_{\mathcal{R}}$ on $\mathcal{L}$-valued powerset algebra $\mathcal{L}^{P}$ of $P$.
Definition 5 ([11]).
Let $(P,\mathcal{R})$ be a Kripke frame and $f\in\mathcal{L}^{P}$. Then $\Box_{\mathcal{R}}f:P\to\mathcal{L}$ is defined by $(\Box_{\mathcal{R}}f)(x)=\bigwedge\{f(y):y\in\mathcal{R}[x]\}$.
Definition 6 ([12]).
Let $\mathcal{A}$ be an object in $\mathcal{MA}_{\mathcal{L}}$. A binary relation $\mathcal{R}_{\Box}$ on $HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ is defined as follows:
$\psi\mathcal{R}_{\Box}\phi\iff\forall\ell\in\mathcal{L},\forall a\in\mathcal{A},\psi(\Box a)\geq\ell\Rightarrow\phi(a)\geq\ell$.
A $\mathcal{L}$-valued map $\mathcal{D}:HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})\times\mathcal{A}\to\mathcal{L}$ is defined by $\mathcal{D}(\psi,a)=\psi(a),\psi\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$.
Lemma 4 ([12]).
The $\mathcal{L}$-valued canonical model $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\mathcal{R}_{\Box},\mathcal{D})$ of $\mathcal{A}$ is a $\mathcal{L}$-valued Kripke model. Then, $\mathcal{D}(\psi,\Box a)=\psi(\Box a)=\bigwedge\{\phi(a):\phi\in\mathcal{R}_{\Box}[\psi]\}$.
2.3 Bitopological duality for Fitting’s Heyting-valued logic
We will introduce the key ideas and findings from the bitopological duality theory for Fitting’s Heyting-valued logic. We refer to [14] for a more thorough explanation of the bitopological duality for Fitting’s Heyting-valued logic. Let $\mathfrak{S}_{\mathcal{L}}$ denote the collection of subalgebras of $\mathcal{L}$. For a pairwise Boolean space $\mathcal{B}$, $\Lambda_{\mathcal{B}}$ denotes the collection of pairwise closed subspaces of $\mathcal{B}$. It is shown in [15] that a pairwise closed subset of a pairwise compact space is also pairwise compact. Hence, each member of $\Lambda_{\mathcal{B}}$ is a pairwise Boolean space. A finite distributive lattice $\mathcal{L}$ endowed with unary operation $T_{\ell}(\ell\in\mathcal{L})$ forms a semi-primal algebra. We have expanded the theory of natural duality [16] by creating a bitopological duality for $\mathcal{L}$-$\mathcal{VL}$-algebras [14].
We now recall the category $PBS_{\mathcal{L}}$ from [14].
Definition 7 ([14]).
The category $PBS_{\mathcal{L}}$ is defined as follows:
  1. (1)
[…]
    1. (i)
      $\alpha_{\mathcal{B}}(\mathcal{L})=\mathcal{B}$;
    2. (ii)
      if $\mathcal{L}_{1}=\mathcal{L}_{2}\wedge\mathcal{L}_{3}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}\in\mathcal{L})$, then $\alpha_{\mathcal{B}}(\mathcal{L}_{1})=\alpha_{\mathcal{B}}(\mathcal{L}_{2})\cap\alpha_{\mathcal{B}}(\mathcal{L}_{3})$.
  2. (2)
[…]
Note 2.
  1. (1)
    The bitopological space $(\mathcal{L},\tau,\tau)$, where $\tau$ is the discrete topology on $\mathcal{L}$, is a pairwise Boolean space. Hence, $(\mathcal{L},\tau,\tau,\alpha_{\mathcal{L}})$, where $\alpha_{\mathcal{L}}$ is a mapping from $\mathfrak{S}_{\mathcal{L}}$ to $\Lambda_{\mathcal{L}}$ that is defined by $\alpha_{\mathcal{L}}(\mathcal{L}^{\prime})=\mathcal{L}^{\prime}$, is an object in $PBS_{\mathcal{L}}$.
  2. (2)
    For an object $\mathcal{A}$ in $\mathcal{VA}_{\mathcal{L}}$ , consider a bitopological space $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2})$, where the topologies $\tau_{1}$ and $\tau_{2}$ are generated by the bases $B^{\tau_{1}}=\{\langle a\rangle:a\in\mathcal{A}\}$, where $\langle a\rangle=\{h\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}):h(a)=1\}$, and $B^{\tau_{2}}=\{B^{c}:B\in B^{\tau_{1}}\}$, respectively. Here, $B^{c}$ denotes the complement of $B$.
Fact 1 ([14]).
The bitopological space $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2})$ is a pairwise Boolean space.
The duality between the categories $\mathcal{VA}_{\mathcal{L}}$ and $PBS_{\mathcal{L}}$ is obtained via the following functors.
Definition 8 ([14]).
A contravariant functor $\mathfrak{F}:PBS_{\mathcal{L}}\to\mathcal{VA}_{\mathcal{L}}$ is defned as follows:
  1. (i)
    For an object $(\mathcal{B},\alpha_{\mathcal{B}})$ in $PBS_{\mathcal{L}}$, define $\mathfrak{F}(\mathcal{B},\alpha_{\mathcal{B}})=(HOM_{PBS_{\mathcal{L}}}((\mathcal{B},\alpha_{\mathcal{B}}),(\mathcal{L},\alpha_{\mathcal{L}})),\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$, where $\vee,\wedge,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1$ are pointwise operations on the set $HOM_{PBS_{\mathcal{L}}}((\mathcal{B},\alpha_{\mathcal{B}}),(\mathcal{L},\alpha_{\mathcal{L}}))$. The operations 00 and $1$ are regarded as constant functions, with 00 and $1$ being their respective values.
  2. (ii)
    For an arrow $\phi:(\mathcal{B},\alpha_{\mathcal{B}})\to(\mathcal{B}^{\prime},\alpha_{\mathcal{B}^{\prime}})$ in $PBS_{\mathcal{L}}$, define $\mathfrak{F}(\phi):\mathfrak{F}((\mathcal{B}^{\prime},\alpha_{\mathcal{B}^{\prime}}))\to\mathfrak{F}((\mathcal{B},\alpha_{\mathcal{B}}))$ by $\mathfrak{F}(\phi)(\zeta)=\zeta\circ\phi$, where $\zeta\in HOM_{PBS_{\mathcal{L}}}((\mathcal{B}^{\prime},\alpha_{\mathcal{B}^{\prime}}),(\mathcal{L},\alpha_{\mathcal{L}}))$.
Definition 9 ([14]).
A contravariant functor $\mathfrak{G}:\mathcal{VA}_{\mathcal{L}}\to PBS_{\mathcal{L}}$ is defined as follows:
  1. (i)
    $\mathfrak{G}$ acts on an object $\mathcal{A}$ in $\mathcal{VA}_{\mathcal{L}}$ as $\mathfrak{G}(\mathcal{A})=(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2},\alpha_{\mathcal{A}})$, where $\alpha_{\mathcal{A}}$ is a mapping from $\mathfrak{S}_{\mathcal{L}}$ to $\Lambda_{HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})}$ which is defined by $\alpha_{\mathcal{A}}(\mathcal{L}^{*})=HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}^{*})$, $\mathcal{L}^{*}\in\mathfrak{S}_{\mathcal{L}}$.
  2. (ii)
    $\mathfrak{G}$ acts on an arrow $\psi:\mathcal{A}\to\mathcal{A}^{*}$ in $\mathcal{VA}_{\mathcal{L}}$ as follows: $\mathfrak{G}(\psi):\mathfrak{G}(\mathcal{A}^{*})\to\mathfrak{G}(\mathcal{A})$ is defined by $\mathfrak{G}(\psi)(\phi)=\phi\circ\psi$, $\phi\in\mathfrak{G}(\mathcal{A}^{*})$.
In [14], the following duality result is proved for $\mathcal{L}$-$\mathcal{VL}$-algebras:
Theorem 1.
The categories $\mathcal{VA}_{\mathcal{L}}$ and $PBS_{\mathcal{L}}$ are dually equivalent.
3 Bitopological duality for Fitting’s many-valued modal logic
Let $\mathcal{R}$ be a relation on $P$ and $C\subseteq P$. We define
$[\mathcal{R}]C=\{p\in P:\mathcal{R}[p]\subseteq C\}$ and $\langle\mathcal{R}\rangle C=\{p\in P:\mathcal{R}[p]\cap C\neq\emptyset\}$.
3.1 Category
Definition 10.
We define a category $PRBS_{\mathcal{L}}$ as follows:
  1. (1)
    Objects: An object in $PRBS_{\mathcal{L}}$ is a triple $(P,\alpha_{P},\mathcal{R})$ such that $(P,\alpha_{P})$ is an object in $PBS_{\mathcal{L}}$ and $\mathcal{R}$ is a binary relation on $P$ that satisfies the following conditions:
    1. (i)
      for each $p$ in $P$, $\mathcal{R}[p]$ is a pairwise compact subset of $P$;
    2. (ii)
      $\forall\mathcal{C}\in\beta_{1}$, $[\mathcal{R}]\mathcal{C},\langle\mathcal{R}\rangle\mathcal{C}\in\beta_{1}$;
    3. (iii)
      for any $\mathcal{L^{\prime}}\in\mathfrak{S}_{\mathcal{L}}$, if $m\in\alpha_{P}(\mathcal{L^{\prime}})$ then $\mathcal{R}[m]\subseteq\alpha_{P}(\mathcal{L^{\prime}})$.
  2. (2)
    Arrows: An arrow $f:(P,\alpha_{P},\mathcal{R})\to(P^{\prime},\alpha_{P^{\prime}},\mathcal{R}^{\prime})$ in $PRBS_{\mathcal{L}}$ is an arrow in $PBS_{\mathcal{L}}$ which additionally satisfies the following conditions:
    1. (i)
      if $p_{1}\mathcal{R}p_{2}$ then $f(p_{1})\mathcal{R}^{\prime}f(p_{2})$;
    2. (ii)
      if $f(p)\mathcal{R}^{\prime}p^{\prime}$ then $\exists p^{*}\in P$ such that $p\mathcal{R}p^{*}$ and $f(p^{*})=p^{\prime}$.
Note 3.
We see that $[\mathcal{R}]Q^{c}=(\langle\mathcal{R}\rangle Q)^{c}$, and $\langle\mathcal{R}\rangle Q^{c}=([\mathcal{R}]Q)^{c}$. Since $\beta_{2}=\{Q^{c}:Q\in\beta_{1}\}$, hence if the relation $\mathcal{R}$ satisfies condition $(ii)$ that is given in the object part of Definition 10, then $[\mathcal{R}]Q,\langle\mathcal{R}\rangle Q\in\beta_{2}$, $\forall Q\in\beta_{2}$.
To prove that the category $\mathcal{MA}_{\mathcal{L}}$ is dually equivalent to the category $PRBS_{\mathcal{L}}$, we define two functors $\mathcal{F}$ and $\mathcal{G}$ in the next sequel.
3.2 Functors
Definition 11.
We define a functor $\mathcal{G}:\mathcal{MA}_{\mathcal{L}}\to PRBS_{\mathcal{L}}$.
  1. (i)
    For an object $(\mathcal{A},\Box)$ in $\mathcal{MA}_{\mathcal{L}}$, define $\mathcal{G}(\mathcal{A})=(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2},\alpha_{\mathcal{A}},\mathcal{R}_{\Box})$, where $\alpha_{\mathcal{A}}:\mathfrak{S}_{\mathcal{L}}\to\Lambda_{HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})}$ is a mapping which is defined by $\alpha_{\mathcal{A}}(\mathcal{L}_{1})=HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}_{1})$, and $\mathcal{R}_{\Box}$ is a binary relation on $HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ that is described in Definition 6.
  2. (ii)
    $\mathcal{G}$ acts on an arrow $\psi:\mathcal{A}_{1}\to\mathcal{A}_{2}$ in $\mathcal{MA}_{\mathcal{L}}$ as follows:
    Define $\mathcal{G}(\psi):\mathcal{G}(\mathcal{A}_{2})\to\mathcal{G}(\mathcal{A}_{1})$ by $\mathcal{G}(\psi)(\phi)=\phi\circ\psi$, where $\phi\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A}_{2},\mathcal{L})$.
Lemma 5 and Lemma 6 demonstrate the well-definedness of $\mathcal{G}$.
Lemma 5.
For an object $(\mathcal{A},\Box)$ in $\mathcal{MA}_{\mathcal{L}}$, $\mathcal{G}(\mathcal{A})$ is an object in $PRBS_{\mathcal{L}}$.
Proof.
Definition 9 shows that $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2},\alpha_{\mathcal{A}})$ is an object in $PBS_{\mathcal{L}}$. Thus, it suffices to demonstrate $\mathcal{R}_{\Box}$ satisfies the requirements given in the object section of Definition 10. First, we ensure that for any $\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$, $\mathcal{R}_{\Box}[\mathcal{W}]\in\delta_{1}\cup\delta_{2}$. Let $\mathcal{U}\not\in\mathcal{R}_{\Box}[\mathcal{W}]$. Then by Definition 6, there is an element $a\in\mathcal{A}$ such that there is $L_{1}\in\mathcal{L}$, for which $\mathcal{W}(\Box a)\geq L_{1}$ but $\mathcal{U}(a)\not\geq L_{1}$. It follows that $\mathcal{U}\in\langle\neg U_{L_{1}}(a)\rangle\in\tau_{2}$ and $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle\neg U_{L_{1}}(a)\rangle=\emptyset$ i.e., $\langle\neg U_{L_{1}}(a)\rangle\subseteq(\mathcal{R}_{\Box}[\mathcal{W}])^{c}$. Hence, $\mathcal{U}\not\in\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}$, where $\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}$ denotes the closure of $\mathcal{R}_{\Box}[\mathcal{W}]$ in $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{2})$ Equivalently, we have $\overline{\mathcal{R}_{\Box}[\mathcal{W}]}^{\tau_{2}}\subset\mathcal{R}_{\Box}[\mathcal{W}]$. Therefore, $\mathcal{R}_{\Box}[\mathcal{W}]$ is $\tau_{2}$-closed. Since $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{1},\tau_{2})$ is pairwise compact, by Proposition 2, we have $\mathcal{R}_{\Box}[\mathcal{W}]$ is pairwise compact.
Now we verify the condition $(ii)$ in the object part of Definition 10. Since $\{\langle a\rangle:a\in\mathcal{A}\}\in\beta_{1}$ and $\{\langle T_{1}(a)\rightarrow 0\rangle:a\in\mathcal{A}\}\in\beta_{2}$ are the basis for the topologies $\tau_{1}$ and $\tau_{2}$, respectively, so we show that for each $a\in\mathcal{A}$, $\langle\mathcal{R}_{\Box}\rangle(\langle a\rangle)\in\beta_{1}$ and $[\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$. We see that
$\begin{split}\langle\mathcal{R}_{\Box}\rangle\langle a\rangle&=\{\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}):\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle a\rangle\neq\emptyset\}\\ &=([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}\\ &=\{\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}):\mathcal{R}_{\Box}[\mathcal{W}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle\}\end{split}$
We show that $([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$ is $\tau_{1}$-open and $\tau_{2}$-closed. Let $\mathcal{U}\in([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Then $\mathcal{R}_{\Box}[\mathcal{U}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle$. It is easy to see that $\exists$ $\tau_{1}$-open set $\langle\Box T_{1}(a)\rangle$ such that $\mathcal{U}\in\langle\Box T_{1}(a)\rangle$. Let $\mathcal{E}\in\langle\Box T_{1}(a)\rangle$. Then $\mathcal{E}(\Box T_{1}(a))=1$. Using the Kripke condition we have $1=\mathcal{E}(\Box T_{1}(a))=\bigwedge\{\mathcal{U}(T_{1}(a)):\mathcal{E}\mathcal{R}_{\Box}\mathcal{U}\}$. According to Lemma 2, $\mathcal{U}(T_{1}(a))$ is either 00 or $1$. Henceforth, for all $\mathcal{U}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ with $\mathcal{E}\mathcal{R}_{\Box}\mathcal{U}$ we have $\mathcal{U}(T_{1}(a))=1$. As a result, $\mathcal{R}_{\Box}[\mathcal{E}]\not\subset\langle T_{1}(a)\rightarrow 0\rangle$ i.e., $\mathcal{E}\in([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Henceforth, $\mathcal{U}\in\langle\Box T_{1}(a)\rangle\subset([\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$. Therefore, $[\mathcal{R}_{\Box}]\langle T_{1}(a)\rightarrow 0\rangle)^{c}$ is $\tau_{1}$-open i.e., $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle$ is $\tau_{1}$-open.
Let $\mathcal{W}\in(\langle\mathcal{R}_{\Box}\rangle\langle a\rangle)^{c}$. Then $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle a\rangle=\emptyset$. It is easy to see that there is $\tau_{1}$-open set $\langle\Box(T_{1}(a)\rightarrow 0)\rangle$ such that $\mathcal{W}\in\langle\Box(T_{1}(a)\rightarrow 0)\rangle$. Also, by applying the Kripke condition, we have $\langle\Box(T_{1}(a)\rightarrow 0)\rangle\subset(\langle\mathcal{R}_{\Box}\rangle\langle a\rangle)^{c}$. Therefore, $\mathcal{W}\in\langle\Box(T_{1}(a)\rightarrow 0)\rangle\subset(\langle\mathcal{R}_{\Box}\rangle\langle a\rangle)^{c}$. It shows that $(\langle\mathcal{R}_{\Box}\rangle\langle a\rangle)^{c}$ is $\tau_{1}$-open i.e., $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle$ is $\tau_{1}$-closed. It follows from Proposition 2 that $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle$ is pairwise compact. Since the topological space $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}),\tau_{2})$ with basis $\{\langle T_{1}(a)\rightarrow 0\rangle:a\in\mathcal{A}\}$ is a Hausdorff space, so $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle$ is $\tau_{2}$-closed. Hence, $\langle\mathcal{R}_{\Box}\rangle\langle a\rangle\in\beta_{1}$.
Next, we show that $[\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$. We see that
$\begin{split}[\mathcal{R}_{\Box}]\langle a\rangle&=\{\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}):\mathcal{R}_{\Box}[\mathcal{W}]\subseteq\langle a\rangle\}\\ &=(\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle)^{c}\end{split}$
We claim that $\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle=\langle\Box T_{1}(a)\rightarrow 0\rangle$. Let $\mathcal{W}\in\langle\Box T_{1}(a)\rightarrow 0\rangle$. Then $\mathcal{W}(\Box T_{1}(a)\rightarrow 0)=1$. Hence, $\mathcal{W}(\Box T_{1}(a))=0$. Using the Kripke condition, we have, $0=\mathcal{W}(\Box T_{1}(a))=\bigwedge\{\mathcal{U}(T_{1}(a)):\mathcal{W}\mathcal{R}_{\Box}\mathcal{U}\}$. Since $\mathcal{U}(T_{1}(a))=0\text{ or }1$, hence $\exists$ $\mathcal{U}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ with $\mathcal{W}\mathcal{R}_{\Box}\mathcal{U}$ such that $\mathcal{U}(T_{1}(a))=0$. Then $\mathcal{U}\in\langle T_{1}(a)\rightarrow 0\rangle$. Therefore, $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle T_{1}(a)\rightarrow 0\rangle\neq\emptyset$. Thus, $\mathcal{W}\in\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle$. Similarly, by employing the Kripke condition, we can show that if $\mathcal{W}\in\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle$ then $\mathcal{W}\in\langle\Box T_{1}(a)\rightarrow 0\rangle$. Since $\langle\Box T_{1}(a)\rightarrow 0\rangle\in\beta_{2}$, we have $\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle\in\beta_{2}$. As a result, $[\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$.
Finally, we demonstrate that $\mathcal{G}(\mathcal{A})$ meets condition $(iii)$ in the object part of Definition 10. Let $u\in\alpha_{\mathcal{A}}(\mathcal{L}^{\prime})=HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L}^{\prime})$. Suppose $\mathcal{R}_{\Box}[u]\not\subset\alpha_{\mathcal{A}}(\mathcal{L}^{\prime})$. Then $\exists v\in\mathcal{R}_{\Box}[u]$ such that $v\notin\alpha_{\mathcal{A}}(\mathcal{L}^{\prime})$. Hence, $\exists a^{*}\in\mathcal{A}$ such that $v(a^{*})\notin\mathcal{L}^{\prime}$. Let $v(a^{*})=\ell^{*}$. Now for any element $\psi\in\alpha_{\mathcal{A}}(\mathcal{L}^{\prime})$,
$\psi(T_{\ell^{*}}(a^{*})\rightarrow a^{*})=\left\{\begin{array}[]{ll}\ell^{*}&\text{ if }\psi(a^{*})=\ell^{*}\\ 1&\text{ if }\psi(a^{*})\neq\ell^{*}\\ \end{array}\right.$
Using Kripke condition, we have $u(\Box(T_{\ell^{*}}(a^{*})\rightarrow a^{*}))=\bigwedge\{\psi(T_{\ell^{*}}(a^{*})\rightarrow a^{*}):\psi\in\mathcal{R}_{\Box}[u]\}$. This shows that $u(\Box(T_{\ell^{*}}(a^{*})\rightarrow a^{*}))=\ell^{*}\notin\mathcal{L}^{\prime}$. But this contradicts the fact that $u\in\alpha_{\mathcal{A}}(\mathcal{L}^{\prime})$. As a result, $\mathcal{G}(\mathcal{A})$ satisfies condition $(iii)$. ∎
Lemma 6.
Let $(\mathcal{A}_{1},\Box_{1})$, $(\mathcal{A}_{2},\Box_{2})$ be the objects in $\mathcal{MA}_{\mathcal{L}}$ and $\psi:\mathcal{A}_{1}\to\mathcal{A}_{2}$ be an arrow in $\mathcal{MA}_{\mathcal{L}}$. Then, $\mathcal{G}(\psi)$ is an arrow in $PRBS_{\mathcal{L}}$.
Proof.
Here $\mathcal{G}(\psi):\mathcal{G}(\mathcal{A}_{2})\to\mathcal{G}(\mathcal{A}_{1})$ is defined by $\mathcal{G}(\psi)(\phi)=\phi\circ\psi$, $\phi\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A}_{2},\mathcal{L})$. It follows from Definition 9 that $\mathcal{G}(\psi)$ is an arrow in $PBS_{\mathcal{L}}$. Therefore, it is still necessary to demonstrate that $\mathcal{G}(\psi)$ satisfies conditions $(i)$ and $(ii)$ listed in the arrow portion of Definition 10. We first check condition $(i)$. Let $v_{1}\mathcal{R}_{\Box_{2}}v_{2}$, where $v_{1},v_{2}\in\mathcal{G}(\mathcal{A}_{2})$. We are to show that $\mathcal{G}(\psi)(v_{1})\mathcal{R}_{\Box_{1}}\mathcal{G}(\psi)(v_{2})$. Now, if $v_{1}\circ\psi(\Box_{1}a_{1})\geq\ell$ for $a_{1}\in\mathcal{A}_{1}$ and $\ell\in\mathcal{L}$, then we have $v_{1}(\Box_{2}\psi(a_{1}))\geq\ell$. As $v_{1}\mathcal{R}_{\Box_{2}}v_{2}$, so we get $v_{2}(\psi(a_{1}))\geq\ell$. Hence, $\mathcal{G}(\psi)(v_{1})\mathcal{R}_{\Box_{1}}\mathcal{G}(\psi)(v_{2})$. We then check condition $(ii)$, which is mentioned in the arrow part of Definition 10. This is equivalent to verifying $\mathcal{R}_{\Box_{1}}[\mathcal{G}(\psi)(v_{1})]=\mathcal{G}(\psi)(\mathcal{R}_{\Box_{2}}[v_{1}])$. Let $\mathcal{W}\in\mathcal{R}_{\Box_{1}}[v_{1}\circ\psi]$, where $\mathcal{W}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A}_{1},\mathcal{L})$. Then $(v_{1}\circ\psi)\mathcal{R}_{\Box_{1}}\mathcal{W}$. Suppose $\mathcal{W}\notin\mathcal{G}(\psi)(\mathcal{R}_{\Box_{2}}[v_{1}])$. Then $\mathcal{W}\neq\mathcal{G}(\psi)(v^{*})$, $\forall v^{*}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A}_{2},\mathcal{L})$ such that $v_{1}\mathcal{R}_{\Box_{2}}v^{*}$. As $(HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A}_{1},\mathcal{L}),\tau_{1},\tau_{2})$ is a pairwise Hausdorff space, so we can consider $\mathcal{W}\in\langle a_{1}\rangle$ and $\mathcal{G}(\psi)(v^{*})=v^{*}\circ\psi\in\langle T_{1}(a_{1})\rightarrow 0\rangle$. Since $\mathcal{W}\in\mathcal{R}_{\Box_{1}}[\mathcal{G}(\psi)(v_{1})]$ and $\mathcal{W}(a_{1})=1$, we have $\mathcal{G}(\psi)(v_{1})(\Box_{1}a_{1})=1$ i.e., $(v_{1}\circ\psi)(\Box_{1}a_{1})=1$. Since $v_{1}\mathcal{R}_{\Box_{2}}v^{*}$, we have $\mathcal{G}(\psi)(v_{1})\mathcal{R}_{\Box_{1}}\mathcal{G}(\psi)(v^{*})$ using the condition $(i)$ specified in the arrow part of Definition 10. As $\mathcal{G}(\psi)(v_{1})(\Box_{1}a_{1})=1$, Lemma 4 shows that $\mathcal{G}(\psi)(v^{*})(a_{1})=1$, i.e., $v^{*}\circ\psi\in\langle a_{1}\rangle$. This contradicts the fact that $\mathcal{G}(\psi)(v^{*})\in\langle T_{1}(a_{1})\rightarrow 0\rangle$. Therefore, $\mathcal{R}_{\Box_{1}}[\mathcal{G}(\psi)(v_{1})]\subseteq\mathcal{G}(\psi)(\mathcal{R}_{\Box_{2}}[v_{1}])$. Similarly, we can show the reverse direction. ∎
Definition 12.
We define a functor $\mathcal{F}:PRBS_{\mathcal{L}}\to\mathcal{MA}_{\mathcal{L}}$.
  1. (i)
    Define $\mathcal{F}(P,\alpha_{P},\mathcal{R})=(HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}})),\wedge,\vee,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1,\Box_{\mathcal{R}})$ for an object $(P,\alpha_{P},\mathcal{R})$ in $PRBS_{\mathcal{L}}$. Definition 5 describes the modal operation $\Box_{\mathcal{R}}$. Here $\wedge,\vee,\rightarrow,T_{\ell}$ are pointwise operations defined on the set $HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$.
  2. (ii)
    Let $\psi:(P_{1},\alpha_{P_{1}},\mathcal{R}_{1})\to(P_{2},\alpha_{P_{2}},\mathcal{R}_{2})$ be an arrow in $PRBS_{\mathcal{L}}$. Define $\mathcal{F}(\psi):\mathcal{F}(P_{2},\alpha_{P_{2}},\mathcal{R}_{2})\to\mathcal{F}(P_{1},\alpha_{P_{1}},\mathcal{R}_{1})$ by $\mathcal{F}(\psi)(\phi)=\phi\circ\psi$ for $\phi\in\mathcal{F}(P_{2},\alpha_{P_{2}},\mathcal{R}_{2})$.
Note 4.
If $\psi,\phi:(P,\tau_{1}^{P},\tau_{2}^{P},\alpha_{P})\to(\mathcal{L},\tau,\tau,\alpha_{\mathcal{L}})$ are pairwise continuous maps then $\psi\wedge\phi,\psi\vee\phi,\psi\rightarrow\phi,T_{\ell}(\psi)$ are pairwise continuous maps. Thus, $(HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}})),\wedge,\vee,\rightarrow,T_{\ell}(\ell\in\mathcal{L}),0,1)$ is a $\mathcal{L}$-$\mathcal{VL}$-algebra.
Lemmas 7 and 8 demonstrate that the functor $\mathcal{F}$ is well-defined.
Lemma 7.
Let $(P,\alpha_{P},\mathcal{R})$ be an object in $PRBS_{\mathcal{L}}$. Then, $\mathcal{F}(P,\alpha_{P},\mathcal{R})$ is an object in $\mathcal{MA}_{\mathcal{L}}$.
Proof.
It is clear from Definition 8 that $\mathcal{F}(P,\alpha_{P})$ is an object in $\mathcal{VA}_{\mathcal{L}}$. We shall now make sure that the modal operation $\Box_{\mathcal{R}}$ on $\mathcal{F}(P,\alpha_{P},\mathcal{R})$ is well-defined. Let $\eta\in\mathcal{F}(P,\alpha_{P},\mathcal{R})$. We then verify $\Box_{\mathcal{R}}\eta\in\mathcal{F}(P,\alpha_{P},\mathcal{R})$. For any $\ell\in\mathcal{L}$,
$\begin{split}(\Box_{\mathcal{R}}\eta)^{-1}(\{\ell\})&=\{p\in P:\bigwedge\{\eta(p^{\prime}):p^{\prime}\in\mathcal{R}[p]=\ell\}\\ &=\langle\mathcal{R}\rangle((T_{\ell}(\eta))^{-1}(\{1\}))\cap(\langle\mathcal{R}\rangle((U_{\ell}(\eta))^{-1}(\{0\})))^{c}\end{split}$
As both $T_{\ell}(\eta)$ and $U_{\ell}(\eta)$ are pairwise continuous maps, henceforth $(T_{\ell}(\eta))^{-1}(\{1\})\in\beta_{1}^{P}\cap\beta_{2}^{P}$ and $(U_{\ell}(\eta))^{-1}(\{0\})\in\beta_{1}^{P}\cap\beta_{2}^{P}$, where $\beta_{1}^{P}=\tau_{1}^{P}\cap\delta_{2}^{P}$ and $\beta_{2}^{P}=\tau_{2}^{P}\cap\delta_{1}^{P}$. Therefore, $(\Box_{\mathcal{R}}\eta)^{-1}(\{\ell\})\in\tau_{1}^{P}$. Also, $(\Box_{\mathcal{R}}\eta)^{-1}(\{\ell\})\in\tau_{2}^{P}$. As a result, $\Box_{\mathcal{R}}\eta$ is a pairwise continuous map from $P$ to $\mathcal{L}$. Furthermore, by applying condition $(iii)$ that is stated in the object part of Definition 10, we see that for any subalgebra $\mathcal{M}\in\mathfrak{S}_{\mathcal{L}}$, and if $m\in\alpha_{\mathcal{L}}(\mathcal{M})$ then $(\Box_{\mathcal{R}}\eta)(m)=\bigwedge\{\eta(m^{\prime}):m^{\prime}\in\mathcal{R}[m]\}\in\alpha_{\mathcal{L}}(\mathcal{M})$. Thus $\Box_{\mathcal{R}}\eta$ is a subspace preserving map. Hence, $\Box_{\mathcal{R}}\eta\in\mathcal{F}(P,\alpha_{P},\mathcal{R})$. ∎
Lemma 8.
Let $\psi:(P_{1},\alpha_{P_{1}},\mathcal{R}_{1})\to(P_{2},\alpha_{P_{2}},\mathcal{R}_{2})$ be an arrow in $PRBS_{\mathcal{L}}$. Then, $\mathcal{F}(\psi)$ is an arrow in $\mathcal{MA}_{\mathcal{L}}$.
Proof.
According to Definition 8, $\mathcal{F}(\psi)$ is an arrow in $\mathcal{VA}_{\mathcal{L}}$. Therefore, it is sufficient to demonstrate that $\mathcal{F}(\psi)(\Box_{\mathcal{R}_{2}}\phi_{2})=\Box_{\mathcal{R}_{1}}(\mathcal{F}(\psi)\phi_{2})$, where $\phi_{2}\in HOM_{{PBS}_{\mathcal{L}}}((P_{2},\alpha_{P_{2}}),(\mathcal{L},\alpha_{\mathcal{L}}))$. For any $p_{1}\in P_{1}$, we have $\mathcal{F}(\psi)(\Box_{\mathcal{R}_{2}}\phi_{2})(p_{1})=\Box_{\mathcal{R}_{2}}\phi_{2}\circ\psi(p_{1})=\bigwedge\{\phi_{2}(p_{2}):p_{2}\in\mathcal{R}_{2}[\psi(p_{1})]\}$, and $\Box_{\mathcal{R}_{1}}(\mathcal{F}(\psi)\phi_{2})(p_{1})=\Box_{\mathcal{R}_{1}}(\phi_{2}\circ\psi)(p_{1})=\bigwedge\{\phi_{2}\circ\psi(p):p\in\mathcal{R}_{1}[p_{1}]\}$. As $\psi$ satisfies conditions $(i)$ and $(ii)$ listed in item 2 of Definition 10, it is easy to show that $(\mathcal{F}(\psi)(\Box_{\mathcal{R}_{2}}\phi_{2}))(p_{1})\leq\Box_{\mathcal{R}_{1}}(\mathcal{F}(\psi)\phi_{2})(p_{1})$ and $\Box_{\mathcal{R}_{1}}(\mathcal{F}(\psi)\phi_{2})(p_{1})\leq(\mathcal{F}(\psi)(\Box_{\mathcal{R}_{2}}\phi_{2}))(p_{1})$. As a result, $\mathcal{F}(\psi)(\Box_{\mathcal{R}_{2}}\phi_{2})=\Box_{\mathcal{R}_{1}}(\mathcal{F}(\psi)\phi_{2})$. ∎
3.3 Bitopological Duality for $\mathcal{L}$$-\mathcal{ML}$ Algebras
In this subsection, we develop bitopological duality for algebras of Fitting’s Heyting-valued modal logic.
Theorem 2.
Let $\mathcal{A}$ be a $\mathcal{L}$$-\mathcal{ML}$ algebra. Then $\mathcal{A}$ is isomorphic to $\mathcal{F}\circ\mathcal{G}(\mathcal{A})$ in $\mathcal{MA}_{\mathcal{L}}$.
Proof.
We define $\gamma^{\mathcal{A}}:\mathcal{A}\to\mathcal{F}\circ\mathcal{G}(\mathcal{A})$ by $\gamma^{\mathcal{A}}(a)(g)=g(a)$, where $a\in\mathcal{A}$ and $g\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$. It is known from Theorem 1 that $\gamma^{\mathcal{A}}$ is an isomorphism in the category $\mathcal{VA}_{\mathcal{L}}$. The only thing left to prove is that $\gamma^{\mathcal{A}}$ preserves the modal operation $\Box$, i.e., $\gamma^{\mathcal{A}}(\Box a)=\Box_{\mathcal{R}_{\Box}}\gamma^{\mathcal{A}}(a)$, $a\in\mathcal{A}$. Let $g\in\mathcal{G}(\mathcal{A})$. Then
[…]
Hence the result follows. ∎
Theorem 3.
Consider an object $(P,\alpha_{P},\mathcal{R})$ in $PRBS_{\mathcal{L}}$. Then, $(P,\alpha_{P},\mathcal{R})$ is isomorphic to $\mathcal{G}\circ\mathcal{F}(P,\alpha_{P},\mathcal{R})$ in the category $PRBS_{\mathcal{L}}$.
Proof.
Define $\zeta_{(P,\alpha_{P},\mathcal{R})}:(P,\alpha_{P},\mathcal{R})\to\mathcal{G}\circ\mathcal{F}(P,\alpha_{P},\mathcal{R})$ by $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\psi)=\psi(p)$, where $p\in P$ and $\psi\in HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Theorem 1 shows that $\zeta_{(P,\alpha_{P},\mathcal{R})}$ is a bi-homeomorphism in the category $PBS_{\mathcal{L}}$. We show that $\zeta_{(P,\alpha_{P},\mathcal{R})}$ and $\zeta^{-1}_{(P,\alpha_{P},\mathcal{R})}$ satisfy the conditions given in item 2 of Definition 10. We claim that for any $p,p^{\prime}\in P$, $p^{\prime}\in\mathcal{R}[p]\iff\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. Let $p^{\prime}\in\mathcal{R}[p]$. Suppose $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}\psi)\geq\ell$, where $\ell\in\mathcal{L}$ and $\psi\in HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Then $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}\psi)=(\Box_{\mathcal{R}}\psi)(p)=\bigwedge\{\psi(p^{*}):p*\in\mathcal{R}[p]\}$. Since $p^{\prime}\in\mathcal{R}[p]$ and $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}\psi)\geq\ell$, we have $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})(\psi)\geq\ell$. Hence, $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)\mathcal{R}_{\Box_{\mathcal{R}}}\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})$, i.e., $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. Now we verify if $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$ then $p^{\prime}\in\mathcal{R}[p]$. We verify its contrapositive statement. Suppose $p^{\prime}\notin\mathcal{R}[p]$. By Definition 10, $\mathcal{R}[p]$ is a pairwise compact subset of pairwise Boolean space $P$. Then It is straightforward to demonstrate that $\mathcal{R}[p]$ is pairwise closed. Therefore we can get a $\tau_{1}^{P}$-basis open set $\mathcal{O}\in\beta_{1}^{P}$ such that $p^{\prime}\in\mathcal{O}$ and $\mathcal{O}\subseteq P-\mathcal{R}[p]$, i.e., $\mathcal{O}\cap\mathcal{R}[p]=\emptyset$. Define a mapping $f:P\to\mathcal{L}$ by
$f(p)=\left\{\begin{array}[]{ll}0&\text{ if $p\in\mathcal{O}$}\\ 1&\text{ if $p\in\mathcal{O}^{c}$}\\ \end{array}\right.$
Then $f$ is a pairwise continuous map from $(P,\tau_{1}^{P},\tau_{2}^{P})$ to $(\mathcal{L},\tau,\tau)$. As a result, it can be shown that $f\in HOM_{PBS_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Now, $\Box_{\mathcal{R}}f(p)=\bigwedge\{f(z):z\in\mathcal{R}[p]\}=1$ and $f(p^{\prime})=0$. Hence, $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}f)=1$ but $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})(f)\neq 1$. Therefore, $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\notin\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. Hence, we have for any $p,p^{\prime}\in P$, $p^{\prime}\in\mathcal{R}[p]\iff\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. As a result, $\zeta_{(P,\alpha_{P},\mathcal{R})}$ and $\zeta^{-1}_{(P,\alpha_{P},\mathcal{R})}$ satisfy conditions $(i)$ and $(ii)$ mentioned in item 2 of Definition 10. ∎
Finally, we obtain the bitopological duality for Fitting’s Heyting-valued modal logic.
Theorem 4.
The categories $\mathcal{MA}_{\mathcal{L}}$ and $PRBS_{\mathcal{L}}$ are dually equivalent.
Proof.
Let $ID_{1}$ and $ID_{2}$ be the identity functors on $\mathcal{MA}_{\mathcal{L}}$ and $PRBS_{\mathcal{L}}$, respectively. This theorem will be proved by defining two natural isomorphisms, $\gamma:ID_{1}\to\mathcal{F}\circ\mathcal{G}$ and $\zeta:ID_{2}\to\mathcal{G}\circ\mathcal{F}$. For an object $\mathcal{A}$ in $\mathcal{MA}_{\mathcal{L}}$ define $\gamma^{\mathcal{A}}:\mathcal{A}\to\mathcal{F}\circ\mathcal{G}(\mathcal{A})$ by $\gamma^{\mathcal{A}}(a)(g)=g(a)$, where $a\in\mathcal{A}$ and $g\in\mathcal{G}(\mathcal{A})$. For an object $(P,\alpha_{P},\mathcal{R})$ in $PRBS_{\mathcal{L}}$ define $\zeta_{(P,\alpha_{P},\mathcal{R})}:(P,\alpha_{P},\mathcal{R})\to\mathcal{G}\circ\mathcal{F}(P,\alpha_{P},\mathcal{R})$ by $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\psi)=\psi(p)$, where $p\in P$ and $\psi\in HOM_{PBS_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Then it can be shown that $\gamma$ and $\zeta$ are natural transformations. According to Theorems 2 and 3, $\gamma$ and $\zeta$ are natural isomorphisms. ∎
In the next section, we develop a coalgebraic duality theorem for algebras of Fitting’s Heyting valued modal logic.
4 Coalgebraic Duality Theorem for Fitting’s multi-valued logic
This section intends to define an endofunctor called as $\mathcal{L}$-biVietoris functor, $V_{\mathcal{L}}^{bi}:PBS_{\mathcal{L}}\to PBS_{\mathcal{L}}$. After that, we exhibit that the category $PRBS_{\mathcal{L}}$ is isomorphic to the category $COALG(V_{\mathcal{L}}^{bi})$ of coalgebras for the endofunctor $V_{\mathcal{L}}^{bi}$.
We refer the reader to [5] for an overview of coalgebras. Now, let us review the concepts of coalgebra and coalgebra morphisms.
Definition 13.
A coalgebra for an endofunctor $\mathfrak{T}:\bf{C}\to\bf{C}$ on a category $\bf{C}$, called $\mathfrak{T}$-coalgebra, is defined by a tuple $(C,T)$, where $C$ is an object in $\bf{C}$ and $T:C\to\mathfrak{T}(C)$ is an arrow in $\bf{C}$.
Definition 14.
Let $(C_{1},T_{1})$ and $(C_{2},T_{2})$ be any two $\mathfrak{T}$-coalgebras. Then $f:(C_{1},T_{1})\to(C_{2},T_{2})$ is said to be a $\mathfrak{T}$-coalgebra morphism if $f:C_{1}\to C_{2}$ is an arrow in $\bf{C}$ which satisfies $T_{2}\circ f=\mathfrak{T}(f)\circ T_{1}$.
$\mathfrak{T}$-coalgebras and $\mathfrak{T}$-coalgebra morphisms form a category, denoted by $COALG(\mathfrak{T})$.
4.1 The structure of the endofunctor $V_{\mathcal{L}}^{bi}$
We define the pairwise Vietoris space as follows:
Definition 15.
Let $(S,\tau_{1}^{S},\tau_{2}^{S})$ be a pairwise topological space and $\mathcal{K}(S)$ the set of all pairwise closed subsets of $S$. We define $\Box U=\{C\in\mathcal{K}(S):C\subseteq U\}$ and $\Diamond U=\{C\in\mathcal{K}(S):C\cap U\neq\emptyset\}$, $U\subseteq S$. Let $\beta_{1}^{S}$ and $\beta_{2}^{S}$ be the basis for the topologies $\tau_{1}^{S}$ and $\tau_{2}^{S}$, respectively. The pairwise Vietoris space $V_{P}(S)$ of the pairwise topological space $(S,\tau_{1}^{S},\tau_{2}^{S})$ is defined as a pairwise topological space $(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$, where $\tau_{1}^{V}$ is the topology on $\mathcal{K}(S)$ generated by subbasis $\{\Box U,\Diamond U:U\in\beta_{1}^{S}\}$ and the topology $\tau_{2}^{V}$ on $\mathcal{K}(S)$ is generated by subbasis $\{\Box U,\Diamond U:U\in\beta_{2}^{S}\}$.
We then show that $V_{P}(S)$ is a pairwise Boolean space whenever $S$ is a pairwise Boolean space.
Lemma 9.
If $(S,\tau_{1}^{S},\tau_{2}^{S})$ is a pairwise Boolean space then $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is pairwise zero-dimensional.
Proof.
We shall show that $\beta_{1}^{V}=\tau_{1}^{V}\cap\delta_{2}^{V}$ is a basis for $\tau_{1}^{V}$, where $\delta_{2}^{V}$ is the set of $\tau_{2}^{V}$-closed sets. Let $\mathcal{O}\in\tau_{1}^{V}$. Then $\mathcal{O}$ can be expressed as $\mathcal{O}=\displaystyle\bigcup_{\lambda\in\Lambda}(\bigcap_{r=1}^{n_{\lambda}}\Box C_{r}\cap\bigcap_{t=1}^{m_{\lambda}}\Diamond C_{t})$, $C_{r},C_{t}\in\beta_{1}^{S}=\tau_{1}^{S}\cap\delta_{2}^{S}$. To demonstrate that $\beta_{1}^{V}$ is a basis for $\tau_{1}^{V}$, we have to prove that $\bigcap_{r=1}^{n_{\lambda}}\Box C_{r}\cap\bigcap_{t=1}^{m_{\lambda}}\Diamond C_{t}\in\beta_{1}^{V}$. Because the finite intersection of the members of $\beta_{1}^{V}$ is again in $\beta_{1}^{V}$, it is sufficient to establish that for $C\in\beta_{1}^{S}$, $\Box C,\Diamond C\in\beta_{1}^{V}$. As $\tau_{1}^{V}$ is the topology generated by the subbasis $\{\Box C,\Diamond C:C\in\beta_{1}^{S}\}$, hence $\Box C,\Diamond C\in\tau_{1}^{V}$. Now we see that $(\Box C)^{c}=\Diamond C^{c}$ and $(\Diamond C)^{c}=\Box C^{c}$. Since $C\in\beta_{1}^{S}$, so $C^{c}\in\beta_{2}^{S}$. As a result, $\Box C,\Diamond C\in\delta_{2}^{V}$. Henceforth, $\Box C,\Diamond C\in\beta_{1}^{V}$. Similarly, it can be shown that $\beta_{2}^{V}=\tau_{2}^{V}\cap\delta_{1}^{V}$, $\delta_{1}^{V}$ is the set of $\tau_{1}^{V}$-closed sets, is a basis for $\tau_{2}^{V}$. ∎
Lemma 10.
If $(S,\tau_{1}^{S},\tau_{2}^{S})$ is a pairwise Boolean space then $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is pairwise Hausdorff.
Proof.
Let $C,C^{\prime}\in\mathcal{K}(S)$ and $C\neq C^{\prime}$. Let $z\in C$ such that $z\neq z^{\prime}$, $\forall z^{\prime}\in C^{\prime}$. For each point $z^{\prime}\in C^{\prime}$, we choose disjoint open sets $U_{z^{\prime}}^{c}\in\beta_{2}^{S}$ and $U_{z^{\prime}}\in\beta_{1}^{S}$ (using the condition that $(S,\tau_{1}^{S},\tau_{2}^{S})$ is pairwise Hausdorff space.) containing points $z^{\prime}$ and $z$, respectively. So the collection $\{U_{z^{\prime}}^{c}:z^{\prime}\in C^{\prime}\}$ is $\tau_{2}^{S}$-open covering of $C^{\prime}$. As $C^{\prime}$ is pairwise compact, so there is a finite collection $\{U_{z^{\prime}_{i}}^{c}:i=1,2,\cdots,n\}$ such that $C^{\prime}\subseteq\bigcup_{i=1}^{n}U_{z^{\prime}_{i}}^{c}$. Let $V^{\prime}=\displaystyle\bigcup_{i=1}^{n}U_{z^{\prime}_{i}}^{c}$ and $U=\displaystyle\bigcap_{i=1}^{n}U_{z^{\prime}_{i}}$. As $z\in C\cap U$, hence $C\cap U\neq\emptyset$. Also, $C^{\prime}\cap U=\emptyset$ because $C^{\prime}\subseteq U^{c}$. It follows that $C\in\Diamond U\in\tau_{1}^{V}$ and $C^{\prime}\notin\Diamond U$ i.e., $C^{\prime}\in(\Diamond U)^{c}=\Box U^{c}\in\tau_{2}^{V}$. So we have two disjoint open sets $\Diamond U\in\tau_{1}^{V}$ and $\Box U^{c}\in\tau_{2}^{V}$ containing $C$ and $C^{\prime}$, respectively. ∎
Lemma 11.
If $(S,\tau_{1}^{S},\tau_{2}^{S})$ is a pairwise Boolean space then $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is pairwise compact.
Proof.
It is known from Proposition 1 that $\{\Box U,\Diamond U:U\in\beta_{1}^{S}\cup\beta_{2}^{S}\}$ is a subbasis for the topology $\tau_{1}^{S}\vee\tau_{2}^{S}$. We shall show that every cover of $\mathcal{K}(S)$ by subbasis-open sets has a finite subcover. Let $\mathcal{K}(S)=\bigcup_{\lambda\in\Lambda}\Box U_{\lambda}\cup\bigcup_{i\in I}\Diamond V_{i}$. Consider $S_{1}=S-\bigcup_{i\in I}V_{i}$. Then $S_{1}$ is a pairwise closed subset of $S$. Hence, $S_{1}\in\mathcal{K}(S)$. Since, $S_{1}\notin\Diamond V_{i}$ for each $i\in I$, so that $S_{1}\in\bigcup_{\lambda\in\Lambda}\Box U_{\lambda}$. Then for some $\lambda^{\prime}\in\Lambda$, $S_{1}\in\Box U_{{\lambda}^{\prime}}$. As a result, $S_{1}\subseteq U_{{\lambda}^{\prime}}$ and hence $S-U_{{\lambda}^{\prime}}\subseteq S-S_{1}=\displaystyle\bigcup_{i\in I}V_{i}$. Then, $S=U_{{\lambda}^{\prime}}\cup\displaystyle\bigcup_{i\in I}V_{i}$. As $S$ is pairwise compact, we have $S=U_{\lambda^{\prime}}\cup\bigcup_{i=1}^{n}V_{i}$. Let $A$ be an arbitrary element of $\mathcal{K}(S)$. If $A\subseteq U_{{\lambda}^{\prime}}$ then $A\in\Box U_{{\lambda}^{\prime}}$ otherwise $A\subseteq\bigcup_{i\in I}V_{i}$ i.e., $A\cap V_{i}\neq\emptyset$ for some $i\in\{1,2,\cdots,n\}$. As a result, $A\in\Box U_{\lambda^{\prime}}\cup\bigcup_{i\in I}\Diamond V_{i}$. Therefore, $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is pairwise compact. ∎
Lemmas 9, 10 and 11 establish the following result:
Theorem 5.
If $(S,\tau_{1}^{S},\tau_{2}^{S})$ is a pairwise Boolean space then $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is also a pairwise Boolean space.
We now construct the $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}$.
Definition 16.
We define a $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}:PBS_{\mathcal{L}}\to PBS_{\mathcal{L}}$ as follows:
  1. (i)
    For an object $(S,\tau_{1}^{S},\tau_{2}^{S},\alpha_{S})$ in $PBS_{\mathcal{L}}$, we define $V_{\mathcal{L}}^{bi}(S,\tau_{1}^{S},\tau_{2}^{S},\alpha_{S})=(V_{P}(S),V_{P}\circ\alpha_{S})$ where $\alpha_{S}$ is a mapping from $\mathfrak{S}_{\mathcal{L}}$ to $\Lambda_{S}$, then for any $\mathcal{L}_{1}\in\mathfrak{S}_{\mathcal{L}}$, $V_{P}\circ\alpha_{S}(\mathcal{L}_{1})$ is the pairwise Vietoris space of a pairwise closed subspace (i.e., pairwise Boolean subspace) $\alpha_{S}(\mathcal{L}_{1})$ of $S$;
  2. (ii)
    For an arrow $f:(S_{1},\tau_{1}^{S_{1}},\tau_{2}^{S_{1}},\alpha_{S_{1}})\to(S_{2},\tau_{1}^{S_{2}},\tau_{2}^{S_{2}},\alpha_{S_{2}})$ in $PBS_{\mathcal{L}}$, $V_{\mathcal{L}}^{bi}(f):(V_{P}(S_{1}),V_{P}\circ\alpha_{S_{1}})\to(V_{P}(S_{2}),V_{P}\circ\alpha_{S_{2}})$ is defined by $V_{\mathcal{L}}^{bi}(f)(K)=f[K]$, where $K\in V_{P}(S_{1})$.
We now show the well-definedness of the functor $V_{\mathcal{L}}^{bi}$.
Lemma 12.
Let $(S,\tau_{1}^{S},\tau_{2}^{S},\alpha_{S})$ be an object in $PBS_{\mathcal{L}}$. Then $V_{\mathcal{L}}^{bi}(S,\tau_{1}^{S},\tau_{2}^{S},\alpha_{S})$ is an object in $PBS_{\mathcal{L}}$.
Proof.
Theorem 5 shows that $V_{P}(S)$ is a pairwise Boolean space. Now we shall show that $V_{P}\circ\alpha_{S}$ is a pairwise closed subspace of $V_{P}(S)$. For $\mathcal{L}_{1}\in\mathfrak{S}_{\mathcal{L}}$, an element of $V_{P}(S)\circ\alpha_{S}(\mathcal{L}_{1})$ is a pairwise compact subset of $\alpha_{S}(\mathcal{L}_{1})$. As $\alpha_{S}(\mathcal{L}_{1})$ is also pairwise compact subspace of $S$, so that an element of $V_{P}\circ\alpha_{S}(\mathcal{L}_{1})$ is a pairwise compact subset of $S$. As a result, $V_{P}\circ\alpha_{S}(\mathcal{L}_{1})$ is a subset of $V_{P}(S)$. For $U\in\beta_{1}^{S}$, we get $\Box U\cap V_{P}\circ\alpha_{S}(\mathcal{L}_{1})=\{C\in V_{P}\circ\alpha_{S}(\mathcal{L}_{1}):C\subset U\}=\Box(U\cap\alpha_{S}(\mathcal{L}_{1}))$ and $\Diamond U\cap V_{P}\circ\alpha_{S}(\mathcal{L}_{1})=\{C\in V_{P}\circ\alpha_{S}(\mathcal{L}_{1}):C\cap U\neq\emptyset\}=\Diamond(U\cap\alpha_{S}(\mathcal{L}_{1}))$. Similarly for $U\in\beta_{2}^{S}$. Hence, $V_{P}\circ\alpha_{S}(\mathcal{L}_{1})$ is a pairwise subspace of $V_{P}(S)$. Since $\alpha_{S}(\mathcal{L}_{1})$ is a pairwise Boolean subspace of $S$, by Theorem 5 we have $V_{P}\circ\alpha_{S}(\mathcal{L}_{1})$ is a pairwise Boolean space. Henceforth, $V_{P}\circ\alpha_{S}(\mathcal{L}_{1})$ is a pairwise closed subspace of $V_{P}(S)$.
Now we show that $V_{P}\circ\alpha_{S}$ satisfies the conditions given in the object part of Definition 7. If $\alpha_{S}(\mathcal{L})=S$ then $V_{P}\circ\alpha_{S}(\mathcal{L})=V_{P}(S)$.
Let $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}\in\mathfrak{S}_{\mathcal{L}}$. If $\mathcal{L}_{1}=\mathcal{L}_{2}\cap\mathcal{L}_{3}$ then we show that $V_{P}(\alpha_{S}(\mathcal{L}_{1}))=V_{P}(\alpha_{S}(\mathcal{L}_{2}))\cap V_{P}(\alpha_{S}(\mathcal{L}_{3}))$. Now $V_{P}(\alpha_{S}(\mathcal{L}_{1}))=V_{P}(\alpha_{S}(\mathcal{L}_{2}\cap\mathcal{L}_{3}))=V_{P}(\alpha_{S}(\mathcal{L}_{2})\cap\alpha_{S}(\mathcal{L}_{3}))$. The element structure of $V_{P}(\alpha_{S}(\mathcal{L}_{2})\cap\alpha_{S}(\mathcal{L}_{3}))$ is of the form $P\cap(\alpha_{S}(\mathcal{L}_{2})\cap\alpha_{S}(\mathcal{L}_{3}))$ and $Q\cap(\alpha_{S}(\mathcal{L}_{2})\cap\alpha_{S}(\mathcal{L}_{3}))$, where $P$ and $Q$ are $\tau_{1}^{S}$-closed set and $\tau_{2}^{S}$-closed set, respectively. The elements of $V_{P}(\alpha_{S}(\mathcal{L}_{2}))\cap V_{P}(\alpha_{S}(\mathcal{L}_{3}))$ are of the form $(P_{1}\cap\alpha_{S}(\mathcal{L}_{2}))\cap(P_{2}\cap\alpha_{S}(\mathcal{L}_{3}))$ and $(Q_{1}\cap\alpha_{S}(\mathcal{L}_{2}))\cap(Q_{2}\cap\alpha_{S}(\mathcal{L}_{3}))$, where $P_{1},P_{2}$ are $\tau_{1}^{S}$-closed and $Q_{1},Q_{2}$ are $\tau_{2}^{S}$-closed. Then it is easy to show that $V_{P}(\alpha_{S}(\mathcal{L}_{2})\cap\alpha_{S}(\mathcal{L}_{3}))\subseteq V_{P}(\alpha_{S}(\mathcal{L}_{2}))\cap V_{P}(\alpha_{S}(\mathcal{L}_{3}))$ and $V_{P}(\alpha_{S}(\mathcal{L}_{2}))\cap V_{P}(\alpha_{S}(\mathcal{L}_{3}))\subseteq V_{P}(\alpha_{S}(\mathcal{L}_{2})\cap\alpha_{S}(\mathcal{L}_{3}))$. ∎
Lemma 13.
Let $\eta:(S_{1},\tau_{1}^{S_{1}},\tau_{2}^{S_{1}},\alpha_{S_{1}})\to(S_{2},\tau_{1}^{S_{2}},\tau_{2}^{S_{2}},\alpha_{S_{2}})$ be an arrow in $PBS_{\mathcal{L}}$. Then $V_{\mathcal{L}}^{bi}(\eta)$ is an arrow in $PBS_{\mathcal{L}}$.
Proof.
Given that $f$ is a pairwise continuous map from a pairwise Boolean space $S_{1}$ to a pairwise Boolean space $S_{2}$. Let $K\in V_{P}(S_{1})$. Then $K$ is a pairwise closed subset of $S_{1}$ and hence $K$ is pairwise compact. Now $V_{\mathcal{L}}^{bi}(\eta)(K)=\eta[K]$ is a pairwise compact subset of $S_{2}$. Since $S_{2}$ is a pairwise Boolean space, $\eta[K]$ is a pairwise closed subset of $S_{2}$. As a result, $V_{\mathcal{L}}^{bi}(\eta)(K)\in V_{P}(S_{2})$. To show that $V_{\mathcal{L}}^{bi}(\eta)$ is pairwise continuous, let $C\in\beta_{1}^{S_{2}}$ and $D\in\beta_{2}^{S_{2}}$. Then
$\begin{split}V_{\mathcal{L}}^{bi}(\eta)^{-1}(\Box C)&=\{K\in V_{P}(S_{1}):V_{\mathcal{L}}^{bi}(\eta)(K)\in\Box C\}\\ &=\{K\in\mathcal{K}(S_{1}):\eta[K]\subseteq C\}\\ &=\{K\in\mathcal{K}(S_{1}):K\subseteq\eta^{-1}(C)\}\\ &=\Box\eta^{-1}(C)\end{split}$
and
$\begin{split}V_{\mathcal{L}}^{bi}(\eta)^{-1}(\Diamond C)&=\{K\in V_{P}(S_{1}):V_{\mathcal{L}}^{bi}(\eta)(K)\in\Diamond C\}\\ &=\{K\in\mathcal{K}(S_{1}):\eta[K]\cap C\neq\emptyset\}\\ &=\{K\in\mathcal{K}(S_{1}):K\cap\eta^{-1}(C)\neq\emptyset\}\\ &=\Diamond\eta^{-1}(C)\end{split}$
Similarly, $V_{\mathcal{L}}^{bi}(\eta)^{-1}(\Box D)=\Box\eta^{-1}(D)$ and $V_{\mathcal{L}}^{bi}(\eta)^{-1}(\Diamond D)=\Diamond\eta^{-1}(D)$. Therefore, $V_{\mathcal{L}}^{bi}(\eta)$ is pairwise continuous. It is still necessary to demonstrate that $V_{\mathcal{L}}^{bi}(\eta)$ is subspace preserving. Let $M\in V_{P}\circ\alpha_{S_{1}}(\mathcal{L}_{1})$, $\mathcal{L}_{1}\in\mathfrak{S}_{\mathcal{L}}$. Then $M\subseteq\alpha_{S_{1}}(\mathcal{L}_{1})$. As $\eta$ is an arrow in $PBS_{\mathcal{L}}$, hence $\eta$ is a subspace preserving map. Thus, $\eta(M)\subseteq\alpha_{S_{2}}(\mathcal{L}_{1})$. It shows that $V_{\mathcal{L}}^{bi}(\eta)(M)\subseteq\alpha_{S_{2}}(\mathcal{L}_{1})$. Thus we have $V_{\mathcal{L}}^{bi}(\eta)(M)\in V_{P}\circ\alpha_{S_{2}}(\mathcal{L}_{1})$. ∎
We introduce two functors $\mathfrak{B}$ and $\mathfrak{C}$ between the categories $PRBS_{\mathcal{L}}$ and $COALG(V_{\mathcal{L}}^{bi})$ to show that these two categories are isomorphic.
Definition 17.
We define a functor $\mathfrak{B}:PRBS_{\mathcal{L}}\to COALG(V_{\mathcal{L}}^{bi})$ as follows:
  1. (i)
    For an object $(S,\alpha_{S},\mathcal{R})$ in $PRBS_{\mathcal{L}}$, define $\mathfrak{B}(S,\alpha_{S},\mathcal{R})=(S,\alpha_{S},\mathcal{R}[-])$, where $\mathcal{R}[-]:(S,\alpha_{S})\to V_{\mathcal{L}}^{bi}(S,\alpha_{S})$ is an arrow in $PBS_{\mathcal{L}}$ defined by $\mathcal{R}[s]=\{p\in S:s\mathcal{R}p\}$, $s\in S$;
  2. (ii)
    For an arrow $f:(S_{1},\alpha_{S_{1}},\mathcal{R}_{1})\to(S_{2},\alpha_{S_{2}},\mathcal{R}_{2})$ in $PRBS_{\mathcal{L}}$, define $\mathfrak{B}(f):(S_{1},\alpha_{S_{1}},\mathcal{R}_{1}[-])\to(S_{2},\alpha_{S_{2}},\mathcal{R}_{2}[-])$ by $\mathfrak{B}(f)=f$.
The well-definedness of the functor $\mathfrak{B}$ is shown by the following two lemmas:
Lemma 14.
Let $(S,\alpha_{S},\mathcal{R})$ be an object in $PRBS_{\mathcal{L}}$. Then $\mathfrak{B}(S,\alpha_{S},\mathcal{R})$ is an object in $COALG(V_{\mathcal{L}}^{bi})$
Proof.
We shall show that $\mathcal{R}[-]:(S,\alpha_{S})\to V_{\mathcal{L}}^{bi}(S,\alpha_{S})$ is an arrow in $PBS_{\mathcal{L}}$. By the conditions given in the object part of Definition 10, we know that for each $s\in S$, $\mathcal{R}[s]$ is pairwise compact subset of $S$. As $S$ is pairwise Boolean space, hence $\mathcal{R}[s]$ is a pairwise closed subset of $S$. Thus $\mathcal{R}[s]\in V_{P}(S)$. Let $U\in\beta_{1}^{S}$. Then
[…]
and
[…]
Similarly, for $U\in\beta_{2}^{S}$, $\mathcal{R}[-]^{-1}(\Box U)=[\mathcal{R}]U\in\beta_{2}^{S}$ and $\mathcal{R}[-]^{-1}(\Diamond U)=\langle\mathcal{R}\rangle U\in\beta_{2}^{S}$. Henceforth, $\mathcal{R}[-]$ is a pairwise continuous map. Now we show that $\mathcal{R}[-]$ is subspace preserving. Let $s\in\alpha_{S}(\mathcal{L}^{\prime})$, $\mathcal{L}^{\prime}\in\mathfrak{S}_{\mathcal{L}}$. It is known from Definition 10 that $\mathcal{R}[s]$ is a pairwise compact subset of $\alpha_{S}(\mathcal{L}^{\prime})$. Since $\alpha_{S}(\mathcal{L}^{\prime})$ is itself a pairwise Boolean space, thus we have $\mathcal{R}[s]\in V_{P}\circ\alpha_{S}(\mathcal{L}^{\prime})$. Therefore, $\mathfrak{B}(S,\alpha_{S},\mathcal{R})$ is a $V_{\mathcal{L}}^{bi}$-coalgebra. ∎
Lemma 15.
Let $f:(S_{1},\alpha_{S_{1}},\mathcal{R}_{1})\to(S_{2},\alpha_{S_{2}},\mathcal{R}_{2})$ be an arrow in $PRBS_{\mathcal{L}}$. Then $\mathfrak{B}(f)$ is an arrow in $COALG(V_{\mathcal{L}}^{bi})$.
Proof.
As $f$ is an arrow in $PRBS_{\mathcal{L}}$, $\mathfrak{B}(f)=f:(S_{1},\alpha_{S_{1}},\mathcal{R}_{1}[-])\to(S_{2},\alpha_{S_{2}},\mathcal{R}_{2}[-])$ is a pairwise continuous map. Now using the conditions mentioned in the arrow part of Definition 10, it is straightforward to verify that $\mathcal{R}_{2}[-]\circ f=V_{\mathcal{L}}^{bi}\circ\mathcal{R}_{1}[-]$. Thus $\mathfrak{B}(f)$ is an arrow in $COALG(V_{\mathcal{L}}^{bi})$. ∎
Definition 18.
We define a functor $\mathfrak{C}:COALG(V_{\mathcal{L}}^{bi})\to PRBS_{\mathcal{L}}$ as follows:
  1. (i)
    For an object $((C,\alpha_{C}),\xi)$ in $COALG(V_{\mathcal{L}}^{bi})$, define $\mathfrak{C}((C,\alpha_{C}),\xi)=(C,\alpha_{C},\mathcal{R}_{\xi})$, where $\mathcal{R}_{\xi}$ is a binary relation on $C$ defined by $d\in\mathcal{R}_{\xi}[c]\iff d\in\xi(c)$, $c,d\in C$;
  2. (ii)
    For an arrow $f:((C_{1},\alpha_{C_{1}}),\xi_{1})\to((C_{2},\alpha_{C_{2}}),\xi_{2})$ in $COALG(V_{\mathcal{L}}^{bi})$, define $\mathfrak{C}(f):(C_{1},\alpha_{C_{1}},\mathcal{R}_{\xi_{1}})\to(C_{2},\alpha_{C_{2}},\mathcal{R}_{\xi_{2}})$ by $\mathfrak{C}(f)=f$.
The well-definedness of the functor $\mathfrak{C}$ is shown by Lemma 16 and Lemma 17.
Lemma 16.
For an object $((C,\alpha_{C}),\xi)$ in $COALG(V_{\mathcal{L}}^{bi})$, $\mathfrak{C}((C,\alpha_{C}),\xi)=(C,\alpha_{C},\mathcal{R}_{\xi})$ is an object in $PRBS_{\mathcal{L}}$.
Proof.
In order to show that $\mathfrak{C}((C,\alpha_{C}),\xi)$ is an object in $PRBS_{\mathcal{L}}$, we must verify that $\mathfrak{C}((C,\alpha_{C}),\xi)$ satisfies the conditions given in the object part of Definition 10. For each $c\in C$, $\mathcal{R}_{\xi}[c]=\xi(c)\in V_{P}(C)$. Hence, $\mathcal{R}_{\xi}[c]$ is a pairwise closed subset of $C$. Thus $\mathcal{R}_{\xi}[c]$ is pairwise compact. Let $D\in\beta_{1}^{C}$. Then
$\begin{split}[\mathcal{R}_{\xi}](D)&=\{c\in C:\mathcal{R}_{\xi}[c]\subseteq D\}\\ &=\{c\in C:\xi(c)\subseteq D\}\\ &=\{c\in C:\xi(c)\in\Box D\}\\ &=\xi^{-1}(\Box D)\in\beta_{1}^{C}\end{split}$
and
$\begin{split}\langle\mathcal{R}_{\xi}\rangle D&=\{c\in C:\mathcal{R}_{\xi}[c]\cap D\neq\emptyset\}\\ &=\{c\in C:\xi(c)\cap D\neq\emptyset\}\\ &=\{c\in C:\xi(c)\in\Diamond D\}\\ &=\xi^{-1}(\Diamond D)\in\beta_{1}^{C}\end{split}$
Finally, let $m\in\alpha_{C}(\mathcal{L^{\prime}})$ for $\mathcal{L^{\prime}}\in\mathfrak{S}_{\mathcal{L}}$. As $\xi$ is a subspace preserving map from $(C,\alpha_{C})$ to $V_{\mathcal{L}}^{bi}(C,\alpha_{C})$, we have $\mathcal{R}_{\xi}[m]=\xi(m)\in V_{P}\circ\alpha_{C}(\mathcal{L^{\prime}})$. Henceforth, $\mathcal{R}_{\xi}[m]\subset\alpha_{C}(\mathcal{L^{\prime}})$. ∎
Lemma 17.
For an arrow $f:((C_{1},\alpha_{C_{1}}),\xi_{1})\to((C_{2},\alpha_{C_{2}}),\xi_{2})$ in $COALG(V_{\mathcal{L}}^{bi})$, $\mathfrak{C}(f):(C_{1},\alpha_{C_{1}},\mathcal{R}_{\xi_{1}})\to(C_{2},\alpha_{C_{2}},\mathcal{R}_{\xi_{2}})$ is an arrow in $PRBS_{\mathcal{L}}$.
Proof.
It is straightforward to prove that $\mathfrak{C}$ is an arrow in $PRBS_{\mathcal{L}}$. ∎
Now we obtain the following result:
Theorem 6.
The categories $PRBS_{\mathcal{L}}$ and $COALG(V_{\mathcal{L}}^{bi})$ are isomorphic.
Proof.
We shall show that the categories $PRBS_{\mathcal{L}}$ and $COALG(V_{\mathcal{L}}^{bi})$ are isomorphic via the functors $\mathfrak{B}$ and $\mathfrak{C}$. Let $(S,\alpha_{S},\mathcal{R})$ be an object in $PRBS_{\mathcal{L}}$. Then $\mathfrak{C}\circ\mathfrak{B}(S,\alpha_{S},\mathcal{R})=\mathfrak{C}(S,\alpha_{S},\mathcal{R}[-])=(S,\alpha_{S},\mathcal{R}_{\mathcal{R}[-]})$. Now $t\in\mathcal{R}_{\mathcal{R}[-]}(s)\iff t\in\mathcal{R}[s]$. Thus, $(S,\alpha_{S},\mathcal{R})=\mathfrak{C}\circ\mathfrak{B}(S,\alpha_{S},\mathcal{R})$. Let $((C,\alpha_{C}),\xi)$ be an object in $COALG(V_{\mathcal{L}}^{bi})$. Then $\mathfrak{B}\circ\mathfrak{C}((C,\alpha_{C}),\xi)=\mathfrak{B}(C,\alpha_{C},\mathcal{R}_{\xi})=((C,\alpha_{C}),\mathcal{R}_{\xi}[-])$. We have $c_{2}\in\mathcal{R}_{\xi}[c_{1}]\iff c_{2}\in\xi(c_{1})$. As a result, $((C,\alpha_{C}),\xi)=\mathfrak{B}\circ\mathfrak{C}((C,\alpha_{C}),\xi)$. It is clear that for an arrow $f$ in $COALG(V_{\mathcal{L}}^{bi})$, $\mathfrak{B}\circ\mathfrak{C}(f)=f$ and for an arrow $f$ in $PRBS_{\mathcal{L}}$, $\mathfrak{C}\circ\mathfrak{B}(f)=f$. ∎
Using Theorems 4 and 6, we arrive at the following duality theorem:
Theorem 7.
The categories $\mathcal{MA}_{\mathcal{L}}$ and $COALG(V_{\mathcal{L}}^{bi})$ are dually equivalent.
Thus the modal semi-primal duality for algebras of Fitting’s Heyting-valued modal logic (for more information, see [12]) can potentially be represented in terms of the coalgebras of $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}$.
Finally, based on the preceding theorems, we can conclude:
Theorem 8.
Fitting’s multi-valued modal logic is sound and complete in the sense of coalgebras of the biVietoris functor $V_{\mathcal{L}}^{bi}$.
The same methodology can be used to build both bitopological and coalgebraic duality for Łukasiewicz $n$-valued modal logic, where $n=\{0,\frac{1}{n-1},\frac{2}{n-2},\cdots,\frac{n-2}{n-1},1\}$. It should be noted that $(n,\vee,\wedge,\star,\circledast,\rightarrow,()^{c},0,1)$ is a semi-primal algebra, where the operations $(\vee,\wedge,\star,\circledast,\rightarrow,()^{c},0,1)$ are defined by
$p\vee q$ $=max(p,q)$
$p\wedge q$ $=min(p,q)$
$p\star q$ $=max(0,p+q-1)$
$p\circledast q$ $=min(1,p+q)$
$p\rightarrow q$ $=min(1,1-p+q)$
$p^{c}$ $=1-p$
and $0,1$ are considered as constants i.e., nullary-operations. Algebras of Łukasiewicz-$n$-valued modal logic can be found in [13].
Thus, it is easy to follow, with sort of adjustment, that bitopological and coalgebraic duality can be developed for the general algebraic structure $\mathbb{ISP_{M}}(\mathcal{L})$, $\mathcal{L}$ is a finite algebra, using the same method as explained in this paper.
As an application of this coalgebraic approach to the bitopological duality, we establish the following:
Corollary 1.
The category $COALG(V_{\mathcal{L}}^{bi})$ has cofree coalgebra structure.
Corollary 2.
A final coalgebra exists for the endofunctor $V_{\mathcal{L}}^{bi}$.
5 Conclusion
We have introduced the category $PRBS_{\mathcal{L}}$ and found a duality for the class of all algebras of a version of Fitting’s Heyting-valued modal logic in bitopological language. This has led to an extension of the natural duality theory for modal algebras. We have demonstrated how the theory of coalgebras can be used to characterise the category $PRBS_{\mathcal{L}}$ and thus obtained a coalgebraic description of the bitopological duality for Fitting’s Heyting-valued modal logic. The Vietoris functor on the category $PBS_{\mathcal{L}}$ has been formally constructed in the present study, and we have ultimately shown that coalgebras for this functor provide sound and complete semantics for Fitting’s multi-valued modal logic. As an application of this coalgebraic duality, we establish the existence of a final coalgebra and cofree coalgebras in the category $COALG(V_{\mathcal{L}}^{bi})$.
In light of our current work, we can suggest some future study directions.
A promising avenue for future research could involve demonstrating how lattice-valued intuitionistic modal logic can be characterized by the coalgebras of an endofunctor $V$ on the category $BES$ of bi-topological Esakia spaces (the concept of bitopological Esakia spaces is discussed in [9]). We are still determining, though, how to describe the relation $\mathcal{R}$ on bitopological Esakia spaces in terms of coalgebras of the functor $V$; this seems to be a topic of inquiry at present.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
I express my gratitude to Kumar Sankar Ray and Prakash Chandra Mali for their insightful comments and discussions. This research received no external funding.